OpenStudy (anonymous):

ALGEBRA 1 >.< ughhh !

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y = one thirdx - 3. y = 3x - 14 y = one thirdx + 2 y = -3x + 22 y = -one thirdx + 8

OpenStudy (anonymous):

if it is parallel, to that line, then i has the same slope

OpenStudy (anonymous):

y-4=1/3=(x-6) ?

OpenStudy (anonymous):

yes. now simplify it.

OpenStudy (anonymous):

y-4=3(x-5) ?

OpenStudy (anonymous):

I got y = one thirdx + 2